# State space representations of the Roesser type for convolutional layers

### v1

@@ 1 Introduction and main result
In order to provide a quick take up of information relevant to the widest audience we present our main findings and a concise introduction on the first page. Generalizations of these results, details, as well as the classification of the work in the literature can be found in the following pages.
@@
$\begin{bmatrix}x_{1}[i_{1}+1,i_{2}]\\ x_{2}[i_{1},i_{2}+1]\\ y[i_{1},i_{2}]\end{bmatrix}=\begin{bmatrix}f_{1}&A_{11}&A_{12}&B_{1}\\ f_{2}&A_{21}&A_{22}&B_{2}\\ g&C_{1}&C_{2}&D\end{bmatrix}\begin{bmatrix}1\\ x_{1}[i_{1},i_{2}]\\ x_{2}[i_{1},i_{2}]\\ u[i_{1},i_{2}]\end{bmatrix},$ (12)
where the matrices $f_{1},\ldots,D$ are given by
$\begin{split}&\left[\begin{array}[]{c|c}A_{12}&B_{1}\\ \hline\cr C_{2}&D\end{array}\right]=\left[\begin{array}[]{ccc|c}K[r_{1},r_{2}]&\cdots&K[r_{1},1]&K[r_{1},0]\\ \vdots&\ddots&\vdots&\vdots\\ K[1,r_{2}]&\cdots&K[1,1]&K[1,0]\\ \hline\cr K[0,r_{2}]&\cdots&K[0,1]&K[0,0]\\ \end{array}\right],\\ &\left[\begin{array}[]{c}A_{11}\\ \hline\cr C_{1}\end{array}\right]=\left[\begin{array}[]{cc}0&0\\ I&0\\ \hline\cr 0&I\end{array}\right],~{}\left[\begin{array}[]{c|c}A_{22}&B_{2}\end{array}\right]=\left[\begin{array}[]{cc|c}0&I&0\\ 0&0&I\end{array}\right],\\ &A_{21}=0,\quad f_{1}=0,\quad f_{2}=0,\quad g=b.\end{split}$ (13)
The state signals $(x_{1}[i_{1},i_{2}])_{i_{1},i_{2}\in\mathbb{N}_{0}}$ with $x_{1}[i_{1},i_{2}]\in\mathbb{R}^{n_{1}}$, $n_{1}=c_{\mathrm{out}}r_{1}$, and $(x_{2}[i_{1},i_{2}])_{i_{1},i_{2}\in\mathbb{N}_{0}}$ with $x_{2}[i_{1},i_{2}]\in\mathbb{R}^{n_{2}}$, $n_{2}=c_{\mathrm{in}}r_{2}$ are given inductively by (1) with $x_{1}[i_{1},i_{2}]=0$, $x_{2}[i_{1},i_{2}]=0$ for $[i_{1},i_{2}]\in(\{0\}\times\mathbb{N}_{0})\cup(\mathbb{N}_{0}\times\{0\})$.
Proofs for Theorem 1 and Theorem 2 are provided in Section 4. The realization (13) has essentially been proposed in (Varoufakis et al., 1987) for the case $c_{\mathrm{in}}=c_{\mathrm{out}}=1$, where (1) falls into the class of all-zero systems. The appeal of (13) is that this realization can be efficiently generated and it is minimal for important special cases.
@@
Assume $c_{\mathrm{in}}=c_{\mathrm{out}}$ and $K[r_{1},r_{2}]$ has full column rank. Then the state dimension $n_{1}+n_{2}=c_{\mathrm{out}}r_{1}+c_{\mathrm{in}}r_{2}$ is minimal among all realizations of convolutions (1).
@@ 2 Related work and background
These days, there exist well-developed and convenient libraries for the training of neural networks but they are still hardly applied in safety critical systems due to a lack of safety and robustness guarantees. To analyze the robustness or reachability of NNs, recently many researchers have been using relaxations of the underlying nonlinear activation functions to render the NN amenable to an analysis using linear matrix inequalities (LMIs) and semidefinite programming (Fazlyab et al., 2020, 2019; Gramlich et al., 2023; Pauli et al., 2023a; Newton and Papachristodoulou, 2022; Hu et al., 2020). Furthermore, there has been a growing interest in the LMI-based design of robust neural networks (Pauli et al., 2021, 2023b; Revay et al., 2020, 2023). In order to include convolutional layers in LMI-based NN analysis and design, one can translate them into sparse and redundant fully-connected layers (Goodfellow et al., 2016; Aquino et al., 2022; Pauli et al., 2022), yet this hinders the scalability of such SDP-based methods (Pauli et al., 2023a).
@@
As Gramlich et al. (2023) described, standard 2-D convolutional layers are finite impulse response (FIR) filters that can be represented as 2-D systems using the Roesser model (Roesser, 1975). While both representations allow for LMI-based NN analysis and design, a state space representation of a convolutional layer is a more compact and non-repetitive formulation of a convolutional layer than the more cumbersome formulation as a fully connected layer. Gramlich et al. (2023) automatically generate such a representation using robust control toolboxes based on symbolic variables (Biannic and Roos, 2016), which is effective yet costly and does not scale to state-of-the-art convolutional layers. Our construction is instant, which gives it an advantage over Gramlich et al. (2023) and makes it convenient to use when analyzing and synthesizing convolutional neural networks based on LMIs. What is more our construction considers general convolutional layers, i.e., we present state space representations for strided and dilated convolutions, discuss N-D convolutions and consider multiple input and multiple output channels.
@@
Minimal state space realizations for 2-D/N-D systems, i.e., realizations of the system with minimal state dimension, can be found, e.g., by inspection using corresponding circuit implementations (Kung et al., 1977; Antoniou, 2002). However, establishing a minimal realization for 2-D and N-D systems is by no means trivial and an open problem since many decades, yet for special system classes constructions of minimal state space representations exist (Kung et al., 1977; Varoufakis et al., 1987). The literature encompasses constructions of state space representations for a large number of system classes (Eising, 1978, 1980; Zak et al., 1986; Venkataraman and Levy, 1994; Galkowski, 2005; Doan et al., 2015), e.g., minimal realizations of all-pole and all-zero systems (Varoufakis et al., 1987). In particular, convolutions are FIR filters that fall into the class of all-zero systems and Varoufakis et al. (1987) present a minimal state space representation for single-input-single-output all-zero systems, which conincides with our construction. We further present realizations for multiple-input-multiple-output all-zero systems that arise naturally in the context of CNNs, wherein the input size corresponds to the number of input channels and the output size corresponds to the number of output channels.
@@
The current paper builds on Gramlich et al. (2023) who first suggested to analyze CNNs using a 2-D systems respresentation. The main contribution of this paper is the construction of a state space representation for 2-D convolutional layers, for which we show that it is minimal for $c_{\mathrm{in}}=c_{\mathrm{out}}$. We further present state space representations for a large class of convolutional layers, including 1-D, 2-D and N-D convolutions, considering stride and dilation. The remainder of the paper is organized as follows. In Section 3, we introduce N-D convolutional layers and state the problem and in Section 4, we provide proofs for our main results stated on the first page of this paper. In Section 5, we then present state space representations for more general convolutions and provide multiple examples to illustrate the concept of our construction. Finally, in Section 6, we conclude the paper.
@@ 3 Problem statement
Our multi-index description of a convolution (14) encompasses N-D convolutions ($d=N$). A 1-D convolution ($d=1$) operates on a 1-D signal, e.g., a time signal, a 2-D convolution ($d=2$) operates on signals with two propagation dimensions, e.g., images, and an N-D convolution considers inputs with even more input dimensions, e.g., 3-D convolutions may be used to process videos or medical images.
The problem considered in this paper is to determine a state-space model of a convolutional layer (14). Theorem 1 deals with the special case $d=2$, and in Section 5, we further provide state space respresentations for extensions of (14).
@@
An affine N-D system $\ell_{2e}^{c_{\mathrm{in}}}(\mathbb{N}_{0}^{d})\to\ell_{2e}^{c_{\mathrm{out}}}(\mathbb{N}_{0}^{d}),(u[\boldsymbol{i}])\mapsto(y[\boldsymbol{i}])$ is described by a Roesser model as
$\begin{bmatrix}x_{1}[\boldsymbol{i}+e_{1}]\\ \vdots\\ x_{d}[\boldsymbol{i}+e_{d}]\\ y[\boldsymbol{i}]\end{bmatrix}=\begin{bmatrix}f_{1}&A_{11}&\cdots&A_{1d}&B_{1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ f_{d}&A_{d1}&\cdots&A_{dd}&B_{d}\\ g&C_{1}&\cdots&C_{d}&D\end{bmatrix}\begin{bmatrix}1\\ x_{1}[\boldsymbol{i}]\\ \vdots\\ x_{d}[\boldsymbol{i}]\\ u[\boldsymbol{i}]\end{bmatrix},$ (28)
where $e_{i}$ denotes the unit vector with $1$ in the $i-th$ position. Here, the collection of matrices $f_{1}$, $f_{2}$, normal-$\ldots$, $C_{d}$, $D$ is called state space representation of the system, $x_{1}[\boldsymbol{i}]\in\mathbb{R}^{n_{1}}$, normal-$\ldots$, $x_{d}[\boldsymbol{i}]\in\mathbb{R}^{n_{d}}$ are the states, $u[\boldsymbol{i}]\in\mathbb{R}^{c_{\mathrm{in}}}$ is the input and $y[\boldsymbol{i}]\in\mathbb{R}^{c_{\mathrm{out}}}$ is the output of the system. We call (28) a linear time-invariant N-D system ($N=d$) if $f_{1}=f_{2}=\ldots=f_{d}=0$ and $g=0$. Otherwise, we call the system affine time-invariant. We consider zero initial conditions, i.e., $x_{1}[\boldsymbol{i}]=0$, …, $x_{d}[\boldsymbol{i}]=0$ for $[\boldsymbol{i}]\in(\{0\}\times\dots\times\{0\}\times\mathbb{N}_{0})\cup\dots\cup(\mathbb{N}_{0}\times\{0\}\times\dots\times\{0\})$.
@@
$\left[\begin{array}[]{c|c|c}\boldsymbol{f}&\boldsymbol{A}&\boldsymbol{B}\\ \hline\cr\boldsymbol{g}&\boldsymbol{C}&\boldsymbol{D}\end{array}\right]:=\left[\begin{array}[]{c|ccc|c}f_{1}&A_{11}&\cdots&A_{1d}&B_{1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ f_{d}&A_{d1}&\cdots&A_{dd}&B_{d}\\ \hline\cr g&C_{1}&\cdots&C_{d}&D\end{array}\right].$
@@
In the following section, we are concerned with the proofs of Theorem 1 and Theorem 2.
@@ 4 Proofs for Theorem 1 and Theorem 2
as
$G(z_{1},z_{2})$
$=\begin{bmatrix}C_{1}&C_{2}\end{bmatrix}\left(\begin{bmatrix}z_{1}I&0\\ 0&z_{2}I\end{bmatrix}-\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{bmatrix}\right)^{-1}\begin{bmatrix}B_{1}\\ B_{2}\end{bmatrix}+D$ (36)
$=\begin{bmatrix}C_{1}&C_{2}\end{bmatrix}\begin{bmatrix}z_{1}I-A_{11}&-A_{12}\\ 0&z_{2}I-A_{22}\end{bmatrix}^{-1}\begin{bmatrix}B_{1}\\ B_{2}\end{bmatrix}+D$ (42)
@@
[Proof of Theorem 2] W.l.o.g. assume $b=0$. The transfer matrix of (1) is given by (29), wherein we interpret $K[r_{1},r_{2}]$ as the leading matrix coefficient of this transfer matrix.
@@
For $|z_{1}|,|z_{2}|>\|A\|$, this transfer function can be represented using the Neuman series as
@@
where the matrix on the right has $r$ copies of $K[r_{1},r_{2}]$ on its diagonal. Since we assumed that $K[r_{1},r_{2}]$ has full column rank, this implies that the matrix on the right has rank $c_{\mathrm{in}}r$. Hence, the codomain of $\begin{bmatrix}A^{r_{1}+r_{2}-1}B&\cdots&AB&B\end{bmatrix}$ must have at least dimension $c_{\mathrm{in}}r$. This implies that $A$ must at least be an $rc_{\mathrm{in}}\times rc_{\mathrm{in}}$ matrix, i.e., the state dimension of the Roesser model must at least be $rc_{\mathrm{in}}$. By $r=r_{1}+r_{2}$ and $c_{\mathrm{in}}=c_{\mathrm{out}}$ we proved of our claim.
5 More general convolutions and examples
In this section, we go through exemplary realizations of convolutions as 2-D systems starting from the simple 1-D convolution case and including also more general convolutions, that is N-D convolutions, strided convolutions and dilated convolutions.

### v2

@@ 1 Introduction and main result
In order to provide quick access to information relevant to the widest audience, we present our main findings and a concise introduction on the first page. Generalizations of these results, details, as well as the classification of the work in the literature can be found in the following pages.
@@
$\begin{bmatrix}x_{1}[i_{1}+1,i_{2}]\\ x_{2}[i_{1},i_{2}+1]\\ y[i_{1},i_{2}]\end{bmatrix}=\begin{bmatrix}0&A_{11}&A_{12}&B_{1}\\ 0&A_{21}&A_{22}&B_{2}\\ g&C_{1}&C_{2}&D\end{bmatrix}\begin{bmatrix}1\\ x_{1}[i_{1},i_{2}]\\ x_{2}[i_{1},i_{2}]\\ u[i_{1},i_{2}]\end{bmatrix},$ (2)
where the matrices $A_{11},\ldots,D$ are given by
$\left[\begin{array}[]{c|c}A_{12}&B_{1}\\ \hline\cr C_{2}&D\end{array}\right]=\left[\begin{array}[]{ccc|c}K[r_{1},r_{2}]&\cdots&K[r_{1},1]&K[r_{1},0]\\ \vdots&\ddots&\vdots&\vdots\\ K[1,r_{2}]&\cdots&K[1,1]&K[1,0]\\ \hline\cr K[0,r_{2}]&\cdots&K[0,1]&K[0,0]\\ \end{array}\right],$ (9)
$\left[\begin{array}[]{c}A_{11}\\ \hline\cr C_{1}\end{array}\right]=\left[\begin{array}[]{cc}0&0\\ I_{d_{\mathrm{out}}}&0\\ \hline\cr 0&I_{c_{\mathrm{out}}}\end{array}\right],~{}\left[\begin{array}[]{c|c}A_{22}&B_{2}\end{array}\right]=\left[\begin{array}[]{cc|c}0&I_{d_{\mathrm{in}}}&0\\ 0&0&I_{c_{\mathrm{in}}}\end{array}\right],$ (18)
$A_{21}=0,~{}g=b,~{}d_{\mathrm{in}}=c_{\mathrm{in}}(r_{2}-1),~{}d_{\mathrm{out}}=c_{\mathrm{out}}(r_{1}-1).$
The state signals $(x_{1}[i_{1},i_{2}])_{i_{1},i_{2}\in\mathbb{N}_{0}}$ with $x_{1}[i_{1},i_{2}]\in\mathbb{R}^{n_{1}}$, $n_{1}=c_{\mathrm{out}}r_{1}$, and $(x_{2}[i_{1},i_{2}])_{i_{1},i_{2}\in\mathbb{N}_{0}}$ with $x_{2}[i_{1},i_{2}]\in\mathbb{R}^{n_{2}}$, $n_{2}=c_{\mathrm{in}}r_{2}$ are given inductively by (1) with $x_{1}[0,i_{2}]=0$ for all $i_{2}\in\mathbb{N}_{0}$, and $x_{2}[i_{1},0]=0$ for all $i_{1}\in\mathbb{N}_{0}$.
The realization (9) has essentially been proposed in (Varoufakis et al., 1987) for the case $c_{\mathrm{in}}=c_{\mathrm{out}}=1$, where (1) falls into the class of all-zero systems. The appeal of (9) is that this realization can be efficiently generated and it is minimal for important special cases.
@@
Assume $c_{\mathrm{in}}=c_{\mathrm{out}}$ and $K[r_{1},r_{2}]$ has full rank. Then the state dimension $n_{1}+n_{2}=c_{\mathrm{out}}r_{1}+c_{\mathrm{in}}r_{2}$ is minimal among all realizations of convolutions (1).
Proofs for Theorems 1 and 2 are provided in Section 4.
@@ 2 Related work and background
Despite their widespread success in various applications that range from time series and image classification to image segmentation and face recognition (Li et al., 2021), CNNs are hardly used in safety-critical systems due to a lack of safety and robustness guarantees. To analyze the robustness or reachability of NNs, many researchers have recently used relaxations of the underlying nonlinear activation functions to render the NN amenable to an analysis using linear matrix inequalities (LMIs) and semidefinite programming (SDP) (Fazlyab et al., 2020, 2019; Gramlich et al., 2023; Pauli et al., 2023a; Newton and Papachristodoulou, 2023; Hu et al., 2020). Furthermore, there has been a growing interest in LMI-based design of robust NNs (Pauli et al., 2021, 2023b; Revay et al., 2020, 2023). In order to include convolutional layers in LMI-based NN analysis and design, one can translate them into sparse and repetitive fully connected layers (Goodfellow et al., 2016; Aquino et al., 2022; Pauli et al., 2022).
@@
As Gramlich et al. (2023) described, standard 2-D convolutional layers are finite impulse response (FIR) filters that can be represented as 2-D systems using the Roesser model (Roesser, 1975). Such a state space representation of a convolutional layer is a non-repetitive and more compact formulation of a convolutional layer than the formulation as a fully connected layer and, therefore, LMI-based methods have better scalability using state space representations (Pauli et al., 2023a). Gramlich et al. (2023) automatically generate such representations using robust control toolboxes based on symbolic variables (Biannic and Roos, 2016), which is effective yet costly and does not scale to state-of-the-art convolutional layers. Our construction is instant, which gives it an advantage over Gramlich et al. (2023) and makes it convenient to use when analyzing and synthesizing CNNs based on LMIs. What is more, state space models have recently gained importance in machine learning (Gu et al., 2022; Gu and Dao, 2024).
@@
Finding minimal state space realizations for 2-D/N-D systems, i.e., realizations of the system with minimal state dimension, is by no means trivial and has been an open problem for decades. The literature encompasses constructions of state space representations for a large number of system classes (Eising, 1978, 1980; Zak et al., 1986; Venkataraman and Levy, 1994; Galkowski, 2005; Doan et al., 2015), e.g., minimal realizations of all-pole and all-zero systems (Varoufakis et al., 1987), that are mostly based on construction by inspection using corresponding circuit implementations (Kung et al., 1977; Antoniou, 2002). In particular, convolutions are FIR filters that fall into the class of all-zero systems and Varoufakis et al. (1987) present a minimal state space representation for single-input-single-output all-zero systems, which conincides with our construction. We further present realizations for multiple-input-multiple-output all-zero systems that arise naturally in the context of CNNs, wherein the input size corresponds to the number of input channels and the output size corresponds to the output channels.
@@
The current paper builds on Gramlich et al. (2023) who first suggested to analyze CNNs using a 2-D systems respresentation. The main contribution of this paper is the construction of a state space representation for 2-D convolutional layers, for which we show that it is minimal for $c_{\mathrm{in}}=c_{\mathrm{out}}$. We further present state space representations for a large class of convolutional layers, including 1-D, 2-D and N-D convolutions, considering stride and dilation, and multiple input and output channels.
The remainder of the paper is organized as follows. In Section 3, we introduce N-D convolutional layers and state the problem and in Section 4, we provide proofs for our main results stated on the first page of this paper. In Section 5, we then present state space representations for more general convolutions and give multiple examples to illustrate the concept of our construction. Finally, in Section 6, we conclude the paper.
@@ 3 Problem statement
Our multi-index description of a convolution (19) encompasses d-D convolutions. A 1-D convolution ($d=1$) operates on a 1-D signal, e.g., a time signal, a 2-D convolution ($d=2$) operates on signals with two propagation dimensions, e.g., images, and an N-D convolution ($d=N$) considers inputs with even more input dimensions, e.g., 3-D convolutions may be used to process videos or medical images.
The problem considered in this paper is to determine a state space model of a convolutional layer (19). Theorem 1 deals with the special case $d=2$, and in Section 5, we provide state space representations for extensions of (19).
@@
A d-D system $\ell_{2e}^{c_{\mathrm{in}}}(\mathbb{N}_{0}^{d})\to\ell_{2e}^{c_{\mathrm{out}}}(\mathbb{N}_{0}^{d}),(u[\boldsymbol{i}])\mapsto(y[\boldsymbol{i}])$ is described by a Roesser model
$\begin{bmatrix}x_{1}[\boldsymbol{i}+e_{1}]\\ \vdots\\ x_{d}[\boldsymbol{i}+e_{d}]\\ y[\boldsymbol{i}]\end{bmatrix}=\begin{bmatrix}0&A_{11}&\cdots&A_{1d}&B_{1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&A_{d1}&\cdots&A_{dd}&B_{d}\\ g&C_{1}&\cdots&C_{d}&D\end{bmatrix}\begin{bmatrix}1\\ x_{1}[\boldsymbol{i}]\\ \vdots\\ x_{d}[\boldsymbol{i}]\\ u[\boldsymbol{i}]\end{bmatrix},$ (20)
where $e_{i}$ denotes the unit vector with $1$ in the $i$-th position. Here, the collection of matrices $A_{11}$, $\ldots$, $C_{d}$, $D$ is called state space representation of the system, $x_{1}[\boldsymbol{i}]\in\mathbb{R}^{n_{1}}$, $\ldots$, $x_{d}[\boldsymbol{i}]\in\mathbb{R}^{n_{d}}$ are the states, $u[\boldsymbol{i}]\in\mathbb{R}^{c_{\mathrm{in}}}$ is the input and $y[\boldsymbol{i}]\in\mathbb{R}^{c_{\mathrm{out}}}$ is the output of the system. We consider zero initial conditions, i.e., $x_{1}[\boldsymbol{i}]=0$ for all $\boldsymbol{i}\in(\{0\}\times\mathbb{N}_{0}\times\dots\times\mathbb{N}_{0})$, …, $x_{d}[\boldsymbol{i}]=0$ for all $\boldsymbol{i}\in(\mathbb{N}_{0}\times\dots\times\mathbb{N}_{0}\times\{0\})$.
@@
$\left[\begin{array}[]{c|c|c}\boldsymbol{f}&\boldsymbol{A}&\boldsymbol{B}\\ \hline\cr\boldsymbol{g}&\boldsymbol{C}&\boldsymbol{D}\end{array}\right]:=\left[\begin{array}[]{c|ccc|c}0&A_{11}&\cdots&A_{1d}&B_{1}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&A_{d1}&\cdots&A_{dd}&B_{d}\\ \hline\cr g&C_{1}&\cdots&C_{d}&D\end{array}\right].$
@@ 4 Proofs for Theorem 1 and Theorem 2
as $G(z_{1},z_{2})=$
$\begin{bmatrix}C_{1}&C_{2}\end{bmatrix}\left(\begin{bmatrix}z_{1}I&0\\ 0&z_{2}I\end{bmatrix}-\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{bmatrix}\right)^{-1}\begin{bmatrix}B_{1}\\ B_{2}\end{bmatrix}+D$
$=\begin{bmatrix}C_{1}&C_{2}\end{bmatrix}\begin{bmatrix}z_{1}I-A_{11}&-A_{12}\\ 0&z_{2}I-A_{22}\end{bmatrix}^{-1}\begin{bmatrix}B_{1}\\ B_{2}\end{bmatrix}+D$
@@
[Proof of Theorem 2] W.l.o.g. assume $b=0$, as it does not affect the minimal state dimension. The transfer matrix of (1) is given by (21), wherein we interpret $K[r_{1},r_{2}]$ as the leading matrix coefficient of this transfer matrix.
@@
For $|z_{1}|,|z_{2}|>\|A\|$, this transfer function can be represented using the Neumann series as
@@
where the matrix on the right has $r$ copies of $K[r_{1},r_{2}]$ on its diagonal. Since we assumed that $K[r_{1},r_{2}]$ has full rank, this implies that the matrix on the right has rank $c_{\mathrm{in}}r$. Hence, the codomain of $\begin{bmatrix}A^{r_{1}+r_{2}-1}B&\cdots&AB&B\end{bmatrix}$ must have at least dimension $c_{\mathrm{in}}r$. This implies that $A$ must at least be an $rc_{\mathrm{in}}\times rc_{\mathrm{in}}$ matrix, i.e., the state dimension of the Roesser model must at least be $rc_{\mathrm{in}}$. By $r=r_{1}+r_{2}$ and $c_{\mathrm{in}}=c_{\mathrm{out}}$ we proved of our claim.
5 General convolutions and examples
In this section, we go through exemplary realizations of convolutions as 2-D systems starting from the simple 1-D convolution case and including also more general convolutions, that is N-D convolutions, strided convolutions, and dilated convolutions.